You plan to purchase a $1 million home in Great Falls, VA. You will make a 20% down payment, and finance the rest with an $800,000 jumbo mortgage from First National Bank of Finance. The current rate offered on this 30 year, fixed rate mortgage is 6% per year, compounded monthly.

a)         What is the monthly payment for this mortgage (ignore any tax and insurance escrow amounts)?

b)         Just after you make the third monthly mortgage payment, you are transferred to your company's subsidiary in Chapel Hill, NC. Consequently, your house is sold. What is the principal balance due on your loan at month 3 (t=3), just after you've made the third monthly mortgage payment?

* Please use financial calculator if possible.
Answer #1

a)
N=12*30
I/Y=6%/12
PV=-800000
FV=0
CPT PMT=$4,796.40


b)
N=3
I/Y=6%/12
PV=-800000
PMT=4796.40
CPT FV=$797,598.83